OpenStudy (anonymous):

the total surface area ofa cylinder, A, is given by the formula A=2πr2+2πrh Given that A=200cm^2 and h=10, find the value of r

OpenStudy (anonymous):

the equation is: \[A= 2\pi r ^{2} + 2\pi rh\]

OpenStudy (unklerhaukus):

can you solve the equation for r?

OpenStudy (jack1):

hint: y = ax^2 + bx + c = 0

OpenStudy (anonymous):

r=Ahr^2/2pi

OpenStudy (anonymous):

ii don;t think that's right though....

OpenStudy (jack1):

hmm... what i would do is make the whole equation = 0 so A = 200 h = 10 it's already in the form of a quadratic (ax^2 + bx + c ) so if A=2πr^2+2πrh 200 = 2πr^2+2πr10 0 = 2πr^2+20πr - 200 y = ax^2 + bx + c now solve using quadratic formula

OpenStudy (jack1):

is that all making sense @n9767 ?

OpenStudy (jack1):

did u need this at all (for reference)?

OpenStudy (anonymous):

oh sorry my computer's really slow! THank uou SOO much!!! i understand now!!

OpenStudy (jack1):

all good, y w, what'd jew get for ur answer (just out of curiosity)?

OpenStudy (anonymous):

i got 2.5 cm

OpenStudy (anonymous):

is it right?

OpenStudy (jack1):

yep, i got 2.54